# Consistent energy-momentum trace couplings of fluids

### v1

@@ 4 Implications
In the following, we will state the main implications of our work. From these implications, one finds that a vast amount of work done in $f(R,T)$ gravity needs to be revisited, or must be considered incorrect. The latter applies to models linear in the trace of a perfect fluid energy-momentum tensor or models of the form $f(R,T)=f^{(1)}(R)+f^{(2)}(T)$. Moreover, we provide a simple argument why models linear in $T$ (or a perfect fluid) should not have been considered in the first place as such models cannot yield new physics.
@@
In [1], Harko et al. introduced $f(R,T)$ gravity in which curvature is coupled to the trace of the energy-momentum tensor of matter for several matter candidates, including a relativistic perfect fluid. We have shown that in such cases, one has to ensure using the proper matter Lagrangian to derive the field equations and equations of motion as resorting to the on-shell Lagrangian does not give the full picture. Thus, we start our investigation of the implications of our results with comparison between the results in [1].
Using the field equation [1, Eq. (11)] together with  [1, Eq. (20) and (21)] one arrives at
@@
$\frac{1}{2}\Bigl(3\upmu\frac{\partial p}{\partial\upmu}-\upmu^{2}\frac{\partial^{2}p}{\partial\upmu^{2}}\Bigr)\,$, (4.3)
is a function of $\upmu$, we can use the inverse equation of state to re-express this as a function of the pressure $p$. This means, for given equation of state $p(\upmu)$ we can construct the function
@@ 4.2 Models of the form $f(R,T)=f^{(1)}(R)+f^{(2)}(T)$
The important question to ask now is how do the modification inspired by considering the full fluid Lagrangians affect $f(R,T)$ gravity models. Let us substitute the ansatz $f(R,T)=f^{(1)}(R)+f^{(2)}(T)$ into the gravitational field equations Eq. (2.20) (a similar argument can be made for Eq. (3.11)) which yields
@@
This is nothing but $f(R)$ gravity with an usually constructed matter source. No background cosmological observation could ever distinguish this model from $f(R)$ gravity. One might argue that perturbation theory can come to the rescue, however, this is not the case. Any perturbations in the quantities $\tilde{\rho}$ and $\tilde{p}$ could be re-expressed in the original quantities, with no new information added.
@@ 4.3 Models linear in $T$ – a heuristic argument
The previous section shows that no model linear in $T$ is distinguishable from $f(R)$ gravity. However, this is indeed expected from first principles, whether one uses Brown’s fluid approach or Schutz’s approach (or any approach for that matter!). That is because the trace of the perfect fluid energy-momentum tensor in the action is given $T\sqrt{-g}=(-\rho+3p)\sqrt{-g}$, which is thus proportional to $\rho\sqrt{g}$ and $p\sqrt{-g}$. But these are the expressions for the on-shell Lagrangian of a perfect fluid regardless of the full fluid Lagrangian expression. Consequently, any model using a linear trace coupling on-shell must be equivalent to a perfect fluid with rescaled pressure or energy-density. These are precisely our results of Eqs. (2.24) and (3.13). Say we start with a model of the form given in Section 4.3 which is linear in $T$. Setting $f(R,T)=f^{(1)}(R)+\chi T$ in Eq. (2.24) yields
@@
where $\chi$ is some proportionality constant. One can check that indeed such a model will merely affect the equation of state and does not contain any new physics. We show this explicitly by working, for example, with $\rho=n^{1+w}$, which is well-known to give the linear equation of state $p=w\rho$. This follows directly from $n\partial\rho/\partial n=(1+w)\rho$ and the definition of pressure. In this case $T=-\rho+3p=(-1+3w)\rho$ and
@@
$p_{\rm eff}$ $=\tilde{w}\rho\qquad\textrm{where}\qquad\tilde{w}:=w-\frac{1}{2\kappa}\chi(3w-1)(1+w)\,.$ (4.10)
regardless of the complexity of the term $\tilde{w}$, all it did is to simply redefine the equation of state parameter with the same proportionality constant. Likewise, following Schutz’s approach, we write $p=\upmu^{1+1/w}$, which allows us to get the linear equation of state as $\rho=p/w$ from $\rho+p=\upmu\partial p/\partial\upmu$. Then
@@ 4.4 Models with linear equation of state
Since an $f(R,T)$ model that is linear in the trace of the energy-momentum only differs from $f(R)$ gravity by the equation of state, we investigate whether there is a relation between the equations of state using the on-shell and off-shell Lagrangian for this case. Let us start by assuming a linear equation of state $p=W\rho$ in the incorrect matter term of the field equations in Eq. (4.1). This yields the an effective energy density given by
@@
This result is equivalent to starting with $p=\tilde{w}\rho$ and setting $w=0$ in the the expression for $\tilde{w}$.
This tool allows us to salvage models with linear trace of energy momentum tensor in $f(R,T)$ gravity by realizing that such models only affect the equation of state which we can readily map to the correct equations.
@@ 4.5 A brief note on cosmological applications
This yields a neat condition on the function $f$ that corresponds to cosmological epochs of accelerated expansion, namely
@@
However, this is completely expected for models of this type, as suggested in Section 4.2. Therefore Eq. (4.20) is nothing but the condition $w_{\rm eff}<-1/3$ rewritten in a way that suggests that models of this form cannot contain new information.
4.6 Non-trivial models
An immediate implication our work is that only models where $R$ and $T$ couple directly, will contain some potentially interesting new properties for fluid models. More mathematically speaking, we require the function $f$ to satisfy $f_{TR}\neq 0$. This follows from the conservation equation Eq. (2.46) which we state again in slightly different form
@@ 4.6 Non-trivial models
where we emphasise that the right-hand side now does depend on the geometry through the Ricci scalar. In other words, $f_{T}$ must be a function of $R$.
One of the most interesting aspects of models of this type is that entropy perturbations could play a role in cosmological perturbation theory. Since $U^{\nu}\nabla^{\mu}T^{\rm eff}_{\mu\nu}=0$ (as shown in Section 2.5), these entropy perturbation would only affect the perpendicular part of the conversation equation. In Schutz’s velocity potentials approach, we also encounter entropy perturbations. However, in this picture (as seen in Section 3.5), the perturbations affect both the parallel and perpendicular directions of the conservation equation.
@@ 5 Conclusion
There does not exist a unique Lagrangian expression for a relativistic perfect fluid, as several expressions lead to the correct fluid equation. However, we have shown that one can reach the same conclusions regardless of the Lagrangian expression, given an appropriate off-shell Lagrangian is considered. The field equations in the two approaches are not identical, as can be seen in Eqs. (2.20) and (3.11). However, they contain the same physical information for a given equation of state. This is in contrast to [4] where it is stated that the field equations are identical. The in-equivalence of the field equations goes back to the choice of the derived quantity. When varying the energy-momentum tensor, as is required when trace couplings are considered, the variation of the derived quantity ($p$ or $\rho$) leads to second-order derivatives in the other quantity, see Eqs. (2.12) or (3.5). The choice of the derived quantity depends on whether one uses Brown or Schutz’s approach and thus leads to the, in fact expected, difference in the field equations.
@@
We have shown this by considering the fluid Lagrangian given by Brown using Lagrange multipliers and also Schutz’s Lagrangian expression using velocity potentials. The former is considered better as all the independent variables are explicitly expressed in the Lagrangian function and each variable is associated to a physical quantity. Yet, the expression using velocity potentials is equally valid and can have the advantage of having a simpler form to work with as the variables are hidden within the velocity function. By checking the conservation equations, we notice that the divergence of energy-momentum tensor does not vanish and indeed the coupling of geometry with the matter trace leads to contributions that do not preserve the conservation equations, most notably the entropy contributions. These conservation equations have also been verified by deriving them both from the field equations and the fluid’s equations of motion, further confirming their validity and internal consistency.
@@
The most significant implication of the findings is that $f(R,T)$ models with linear expressions of the matter trace do not lead to new physics compared to their $f(R)$ counterparts. They only lead to modified equations of state. Thus, studying the physics of non-minimal trace coupling requires the consideration of strong geometry-matter couplings or higher order models. On a positive note, this finding means that, since models liner in the matter trace only affect the equation of state, we can map the results from such models done using the on-shell Lagrangian to the proper equations of state.
@@
It is clear that our results apply to any modified theory of gravity that contains non-minimal couplings with perfect fluid matter. We expect that many results in this direction require a careful reinvestigation.
@@ Acknowledgments
We thank Tiberiu Harko and Francisco Lobo for their useful comments on this preprint. Eissa Al-Nasrallah acknowledges the funding provided by Kuwait University through its graduate student scholarship programme.
@@ A.2 Variation of Brown’s fluid Lagrangian
Substituting the expression for $\delta n$ derived in Eq. (A.2)
@@ A.3 Conservation equations via field equations in Brown’s approach
Next, use $h^{\nu\alpha}\nabla^{\mu}h_{\mu\nu}=U^{\mu}\nabla_{\mu}U^{\alpha}$
@@
Rename and lower one index
@@ A.4 Conservation equations via fluid equations in Brown’s approach
Now apply $h_{\alpha\sigma}=g_{\alpha\sigma}+U_{\alpha}U_{\sigma}$
@@
Now, apply the contraction and combine terms
@@
$=\big(\rho+p\big)U^{\beta}\nabla_{\beta}U_{\sigma}+(\delta^{\beta}_{\sigma}+U_{\sigma}U^{\beta})\nabla_{\beta}p\,.$ (A.13)
@@
We now add terms $\pm U^{\beta}\nabla_{\sigma}(\upmu U_{\beta})$ which allows us to anti-symmetrize the first term
@@
We can find an expression for the first term on the right-hand side. First, consider the relation in Eq. (2.30)
@@
Taking the skew part of the divergence of this equation
@@
where since $U^{\beta}\nabla_{\alpha}U_{\beta}=0$, we get
@@
$=h^{\sigma}_{\alpha}\Big(n\nabla_{\beta}\Big(h^{\beta}_{\sigma}f_{T}\frac{\partial T}{\partial n}\Big)-f_{T}\frac{\partial T}{\partial s}\nabla_{\sigma}s\Big)\,.$ (A.28)
@@ A.5 Conservation equations via fluid equations in Schutz’s approach
Starting from the effective fluid equation
@@
If we evaluate $\nabla^{\mu}U_{\nu}$ we find
@@
where $h^{\nu}_{\sigma}U_{\nu}=0$. We indeed get the result expected!

### v2

@@ 4 IMPLICATIONS
In the following, we will state the main implications of our work. From these implications, one finds that a vast amount of work done in $f(R,T)$ gravity needs to be revisited or must be considered incorrect. The latter applies to models linear in the trace of a perfect fluid energy-momentum tensor or models of the form $f(R,T)=f^{(1)}(R)+f^{(2)}(T)$. Moreover, we provide a simple argument why models linear in $T$ (or a perfect fluid) should not have been considered in the first place as such models cannot yield new physics.
@@
In [5], Harko et al. introduced $f(R,T)$ gravity in which curvature is coupled to the trace of the energy-momentum tensor of matter for several matter candidates, including a relativistic perfect fluid. We have shown that in such cases one has to ensure using the proper matter Lagrangian to derive the field equations and equations of motion as resorting to the on-shell Lagrangian does not give the full picture. Thus, we start our investigation of the implications of our results with comparison between the results in [5].
Using the field equation [see Eq. (11) in [5]] together with Eqs. (20) and (21) in [5] one arrives at
@@
$\frac{1}{2}\Bigl(3\upmu\frac{\partial p}{\partial\upmu}-\upmu^{2}\frac{\partial^{2}p}{\partial\upmu^{2}}\Bigr)$ (4.3)
is a function of $\upmu$, we can use the inverse equation of state to reexpress this as a function of the pressure $p$. This means, for given equation of state $p(\upmu)$ we can construct the function
@@ 4.2 Models of the form $f(R,T)=f^{(1)}(R)+f^{(2)}(T)$
The important question to ask now is how do the modification inspired by considering the full fluid Lagrangians affect $f(R,T)$ gravity models. Let us substitute the ansatz $f(R,T)=f^{(1)}(R)+f^{(2)}(T)$ into the gravitational field equations (2.20) [a similar argument can be made for Eq. (3.11)] which yields
@@
This is nothing but $f(R)$ gravity with an usually constructed matter source; see [15]. No background cosmological observation could ever distinguish this model from $f(R)$ gravity. One might argue that perturbation theory can come to the rescue; however, this is not the case. Any perturbations in the quantities $\tilde{\rho}$ and $\tilde{p}$ could be reexpressed in the original quantities, with no new information added.
@@ 4.3 Models linear in $T$ – A heuristic argument
The previous section shows that no model linear in $T$ is distinguishable from $f(R)$ gravity. However, this is indeed expected from first principles, whether one uses Brown’s fluid approach or Schutz’s approach (or any approach for that matter!). That is because the trace of the perfect fluid energy-momentum tensor in the action is given $T\sqrt{-g}=(-\rho+3p)\sqrt{-g}$, which is thus proportional to $\rho\sqrt{g}$ and $p\sqrt{-g}$. But these are the expressions for the on-shell Lagrangian of a perfect fluid regardless of the full fluid Lagrangian expression. Consequently, any model using a linear trace coupling on shell must be equivalent to a perfect fluid with rescaled pressure or energy density. These are precisely our results of Eqs. (2.24) and (3.13). Say we start with a model of the form given in Sec. 4.2 which is linear in $T$. Setting $f(R,T)=f^{(1)}(R)+\chi T$ in Eq. (2.24) yields
@@
where $\chi$ is some proportionality constant. One can check that indeed such a model will merely affect the equation of state and does not contain any new physics. We show this explicitly by working, for example, with $\rho=n^{1+w}$, which is well known to give the linear equation of state $p=w\rho$. This follows directly from $n\partial\rho/\partial n=(1+w)\rho$ and the definition of pressure. In this case $T=-\rho+3p=(-1+3w)\rho$ and
@@
$p_{\rm eff}$ $=\tilde{w}\rho\,,\qquad\textrm{where}\quad\tilde{w}:=w-\frac{1}{2\kappa}\chi(3w-1)(1+w)\,.$ (4.10)
regardless of the complexity of the term $\tilde{w}$; all it did is to simply redefine the equation of state parameter with the same proportionality constant. Likewise, following Schutz’s approach, we write $p=\upmu^{1+1/w}$, which allows us to get the linear equation of state as $\rho=p/w$ from $\rho+p=\upmu\partial p/\partial\upmu$. Then
@@ 4.4 Models with linear equation of state
Since an $f(R,T)$ model that is linear in the trace of the energy-momentum only differs from $f(R)$ gravity by the equation of state, we investigate whether there is a relation between the equations of state using the on-shell and off-shell Lagrangian for this case. Let us start by assuming a linear equation of state $p=W\rho$ in the incorrect matter term of the field equations in Eq. (4.1). This yields an effective energy density given by
@@
This result is equivalent to starting with $p=\tilde{w}\rho$ and setting $w=0$ in the expression for $\tilde{w}$.
This tool allows us to salvage models with linear trace of energy-momentum tensor in $f(R,T)$ gravity by realizing that such models only affect the equation of state which we can readily map to the correct equations.
@@ 4.5 A brief note on cosmological applications
This yields a neat condition on the function $f$ that corresponds to cosmological epochs of accelerated expansion, namely,
@@
However, this is completely expected for models of this type, as suggested in Sec. 4.2. Therefore Eq. (4.20) is nothing but the condition $w_{\rm eff}<-1/3$ rewritten in a way that suggests that models of this form cannot contain new information.
4.6 Nontrivial models
An immediate implication our work is that only models where $R$ and $T$ couple directly will contain some potentially interesting new properties for fluid models. More mathematically speaking, we require the function $f$ to satisfy $f_{TR}\neq 0$. This follows from the conservation equation (2.46) which we state again in a slightly different form:
@@ 4.6 Nontrivial models
where we emphasize that the right-hand side now does depend on the geometry through the Ricci scalar. In other words, $f_{T}$ must be a function of $R$.
One of the most interesting aspects of models of this type is that entropy perturbations could play a role in cosmological perturbation theory. Since $U^{\nu}\nabla^{\mu}T^{\rm eff}_{\mu\nu}=0$ (as shown in Sec. 2.5), these entropy perturbations would only affect the perpendicular part of the conversation equation. In Schutz’s velocity potentials approach, we also encounter entropy perturbations. However, in this picture (as seen in Sec. 3.5), the perturbations affect both the parallel and perpendicular directions of the conservation equation.
@@ 5 CONCLUSION
There does not exist a unique Lagrangian expression for a relativistic perfect fluid, as several expressions lead to the correct fluid equation. However, we have shown that one can reach the same conclusions regardless of the Lagrangian expression, given an appropriate off-shell Lagrangian is considered. The field equations in the two approaches are not identical, as can be seen in Eqs. (2.20) and (3.11). However, they contain the same physical information for a given equation of state. This is in contrast to [6] where it is stated that the field equations are identical. The inequivalence of the field equations goes back to the choice of the derived quantity. When varying the energy-momentum tensor, as is required when trace couplings are considered, the variation of the derived quantity ($p$ or $\rho$) leads to second-order derivatives in the other quantity; see Eq. (2.12) or (3.5). The choice of the derived quantity depends on whether one uses Brown or Schutz’s approach and thus leads to the, in fact expected, difference in the field equations.
@@
We have shown this by considering the fluid Lagrangian given by Brown using Lagrange multipliers and also Schutz’s Lagrangian expression using velocity potentials. The former is considered better as all the independent variables are explicitly expressed in the Lagrangian function and each variable is associated to a physical quantity. Yet, the expression using velocity potentials is equally valid and can have the advantage of having a simpler form to work with as the variables are hidden within the velocity function. By checking the conservation equations, we notice that the divergence of energy-momentum tensor does not vanish and indeed the coupling of geometry with the matter trace leads to contributions that do not preserve the conservation equations, most notably the entropy contributions. It has been shown that entropy perturbations in cosmological fluids are associated with large-scale perturbations in curvature hypersurfaces in multifluid interactions [16, 17], a topic to be further investigated in future work. The conservation equations have also been verified by deriving them both from the field equations and the fluid’s equations of motion, further confirming their validity and internal consistency. Additionally, we attempted to identify the new contributions with dark matter and dark energy to better understand their possible interaction. The Schutz formulation lends itself quite naturally to yield a dark matter contribution.
@@
The most significant implication of the findings is that $f(R,T)$ models with linear expressions of the matter trace do not lead to new physics compared to their $f(R)$ counterparts. They only lead to modified equations of state. Thus, studying the physics of nonminimal trace coupling requires the consideration of strong geometry-matter couplings or higher-order models. On a positive note, this finding means that, since models liner in the matter trace only affect the equation of state, we can map the results from such models done using the on-shell Lagrangian to the proper equations of state.
@@
Our results apply to any modified theory of gravity that contains nonminimal couplings with perfect fluid matter. We expect that many results in this direction require a careful reinvestigation.
@@ ACKNOWLEDGMENTS
We thank Tiberiu Harko and Francisco Lobo for their useful comments on this preprint. E.A. acknowledges the funding provided by Kuwait University through its graduate student scholarship program.
DATA AVAILABILITY
No data were created or analyzed in this study.
@@ A.2 Variation of Brown’s fluid Lagrangian
Substituting the expression for $\delta n$ derived in Eq. (A.2),
@@ A.3 Conservation equations via field equations in Brown’s approach
Next, use $h^{\nu\alpha}\nabla^{\mu}h_{\mu\nu}=U^{\mu}\nabla_{\mu}U^{\alpha}$:
@@
Rename and lower one index:
@@ A.4 Conservation equations via fluid equations in Brown’s approach
Now apply $h_{\alpha\sigma}=g_{\alpha\sigma}+U_{\alpha}U_{\sigma}$:
@@
Now, apply the contraction and combine terms:
@@
$=\big(\rho+p\big)U^{\beta}\nabla_{\beta}U_{\sigma}+(\delta^{\beta}_{\sigma}+U_{\sigma}U^{\beta})\nabla_{\beta}p\,$, (A.13)
@@
We now add terms $\pm U^{\beta}\nabla_{\sigma}(\upmu U_{\beta})$ which allows us to antisymmetrize the first term:
@@
We can find an expression for the first term on the right-hand side. First, consider the relation in Eq. (2.30):
@@
Taking the skew part of the divergence of this equation,
@@
where, since $U^{\beta}\nabla_{\alpha}U_{\beta}=0$, we get
@@
$=h^{\sigma}_{\alpha}\Big(n\nabla_{\beta}\Big(h^{\beta}_{\sigma}f_{T}\frac{\partial T}{\partial n}\Big)-f_{T}\frac{\partial T}{\partial s}\nabla_{\sigma}s\Big)\,$, (A.28)
@@ A.5 Conservation equations via fluid equations in Schutz’s approach
Starting from the effective fluid equation,
@@
If we evaluate $\nabla^{\mu}U_{\nu}$, we find
@@
where $h^{\nu}_{\sigma}U_{\nu}=0$. We indeed get the result expected.